The monthly pass of a bus costs Rs. 150.Everytime vikas travels on the bus and Rs. 7is deducted from her balance.How many times she has to deposit more money in her bus pass if the minimum balance allowed is Rs. 30 ?

Grade:11

1 Answers

vijay kale
6 years ago
Money required for entire month=7×30=210Rs.Total Money she needs to pay=30+210=240rs.Money she now needs to pay=240-150=90No. Of times she needs to pay=90/7=13th message more